OpenStudy (anonymous):

Jim began a 130 mile bike trip to build up stamina for a triathlon. Unfortunately, his chain broke, so he finished the trip walking. The whole trip took 6 hours. If Jim walks at 3 mph, and rides at 35 mph, how long was he on the bike?

OpenStudy (anonymous):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

maybe @confluxepic could help...?

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

You need two equations again.

OpenStudy (anonymous):

W=walking and R=riding

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

3w +35r = 6?

OpenStudy (anonymous):

3w + 35r would reflect the distance you went

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

so 3x+35r=130

OpenStudy (anonymous):

Good...now what will be the second equation that represents time on the bike?

OpenStudy (anonymous):

so do i need to use the substitution method for these two equations? 3w + 35r = 130 w + r = 6

OpenStudy (anonymous):

You got this....very good

OpenStudy (anonymous):

w = 6-r so 3(6-r) +35 r = 130 18-3r + 35r = 130 18 +32r =130 32r + = 112

OpenStudy (anonymous):

r = 3.5

OpenStudy (anonymous):

so he rode for 3.5 hours and walked for 2.5

OpenStudy (anonymous):

Thanks for the help :) I think I will be able to do it on my own next time!

OpenStudy (anonymous):

Always plug the numbers back in the equation to check your answers..very good and good luck